# Compression fronts from fast radio bursts

### v1

@@ Abstract
When a fast radio burst (FRB) expands from its source through a surrounding tenuous plasma, it strongly heats and compresses the plasma at radii up to $\sim 10^{14}$ cm. The likely central engines of FRBs are magnetars, and their ambient plasma at radii $r\gg 10^{10}$ cm is a magnetized $e^{\pm}$ wind. We formulate equations describing the FRB-plasma interaction, solve them numerically, and provide an approximate analytical picture of the interaction. We find the following: (1) FRBs emitted at $r<r_{\rm stoch}\sim 10^{12}$ cm induce fast stochastic heating and strong compression of the wind, sweeping it like a broom. The outcome of this interaction is determined by the energy losses of the radio wave. We evaluate the parameter space where the FRB survives its interaction with the wind. (2) At radii $r>r_{\rm stoch}$, the FRB induces regular heating to the Lorentz factor $\sim a_{0}$, where $a_{0}\propto r^{-1}$ is the wave strength parameter. At $r>r_{\star}\sim 10^{13}\,$cm, the FRB drives a quasisteady compression wave in the wind, with compression factor $C_{\star}\approx 1+a_{0}^{2}$. Both characteristic radii $r_{\rm stoch}$ and $r_{\star}$ scale with FRB luminosity $L$ as $L^{1/3}$. FRBs avoid damping if they are released into the wind medium outside $r_{\rm damp}\sim 10^{11}$ cm.
@@
1 Introduction
Observations of fast radio bursts (FRBs) demonstrate the existence of ultrastrong electromagnetic waves in the Universe. They have characteristic frequencies $\nu=\omega/2\pi\sim 1$ GHz and durations $T\sim 1$ ms (Petroff et al., 2019), and their dimensionless strength parameter is [^1] Throughout the paper we use notation $Q_{n}$ for a quantity $Q$ normalized to $10^{n}$ in CGS units, e.g. $\nu_{9}=\nu/10^{9}{\rm Hz}$.
@@ 1 Introduction
where $L=cr^{2}\langle E_{\rm wave}^{2}\rangle$ is the isotropic equivalent of wave luminosity; $\langle E_{\rm wave}^{2}\rangle^{1/2}$ is the root-mean-square of the wave electric field when the wave expands to a radius $r$.
FRBs have to propagate through the ambient plasma surrounding the source. The likely central engines of FRBs are neutron stars, which are surrounded by electron-positron plasma. The plasma is strongly magnetized:
@@
where the magnetic field $B$ and density $\rho$ are measured in the plasma rest frame. Near the neutron star, the plasma is locked in the closed magnetosphere, rotating with a period $P\sim 1-10$ s. Outside the light cylinder $R_{\rm LC}=cP/2\pi\approx 5\times 10^{9}(P/1{\rm\,s})$ cm, a relativistic wind flows from the magnetosphere, powered by the star’s rotation. Both the magnetosphere and the wind have $\sigma\gg 1$ (see Cerutti & Beloborodov (2017) for a review).
The plasma around neutron stars is tenuous, however it can have a strong effect on FRB propagation. The radio wave with $a\gg 1$ accelerates plasma particles to high energies, and this process can damp the wave, in particular if the energized plasma experiences radiative losses. Strong radiative damping has been demonstrated for FRBs emitted near a magnetar and propagating through its equilibrium outer magnetosphere at radii $r\sim 10^{8}-10^{9}$ cm (Beloborodov, 2021, 2024).
@@
The present paper revisits the FRB-wind interaction. The problem is formulated in section 2 and approached in two steps:
First, we use test-particle simulations to demonstrate two types of heating, regular and stochastic, of a plasma exposed to (and compressed by) an FRB (section 3). When the plasma magnetic field is increased by compression, regular oscillations of particles in the radio wave are changed to stochastic resonant heating. This change boosts the heating and makes it irreversible.
@@
Then, we examine how heating by the FRB results in the bulk acceleration and compression of the plasma. Heating determines acceleration because energy $\Delta{\cal E}$ received from the radio wave always comes with momentum $\Delta{\cal E}/c$ (the wave loses momentum $\Delta{\cal E}/c$). Plasma is accelerated and compressed on timescales much longer than the Larmor period and the wave oscillation period. [^2] In simplest cases, the effect may be thought of as a ponderomotive force exerted by the wave. In the presence of strong radiative losses it is more appropriate to view it as a scattering process (B22). In any case, it is described from first principles by solving for the particle motion in the wave and finding $\Delta{\cal E}/c$. Therefore, it behaves as a fluid governed by magnetohydrodynamics (MHD). In section 4, we formulate MHD equations for the plasma bulk motion inside a propagating radio wave packet. These equations determine plasma compression by FRBs.
@@
Before considering FRBs expanding through a radial wind, we examine the compression front in a plane geometry with a uniform background. First, we show that the problem has a simple steady-state solution (section 5). Its compression factor $C_{\star}$ is directly related to the particle energy gained from the radio wave and becomes huge for waves with strength parameter $a\gg 1$. Remarkably, $C_{\star}$ is independent of the plasma magnetization $\sigma$. This fact casts some doubt on the relevance of the steady-state solution; in particular, it clearly becomes irrelevant in the vacuum limit ($\sigma\rightarrow\infty$) where the radio wave should have no compression effect on a background magnetic field. A complete picture is revealed only by the full time-dependent solution (section 6). It shows that the MHD flow pattern in the propagating wave packet is gradually attracted to the steady state, and this relaxation takes a long time $t_{\star}$ proportional to $\sigma$. For propagation times $t<t_{\star}$, the steady state with large $C_{\star}$ is not reached; we evaluate the actual compression factor $C\ll C_{\star}$ in this situation.
@@
These results allow one to investigate the FRB interaction with a magnetar wind in a broad range of radii. We describe the wind in section 7, and then investigate the compression front induced by the FRB in the wind. We identify two zones, with regular and stochastic heating, and find the transition radius $r_{\rm stoch}$ between them. Compression fronts from FRBs emitted in each zone are described in sections 8 and 9. FRB damping by its interaction with the wind is investigated in section 10. Our conclusions are summarized in section 11.
2 Problem formulation
2.1 The role of magnetic field
@@
For an unmagnetized plasma, particle acceleration in a strong electromagnetic wave is a textbook problem (Landau & Lifshitz, 1975) well known in the context of laser experiments (Bulanov et al., 2015). It was also discussed in the context of early pulsar models (Gunn & Ostriker, 1971). The wave accelerates initially static particles to Lorentz factors $\gamma=1+a^{2}/2$, and their average velocity is aligned with the wave propagation direction. This average motion gives the plasma the bulk Lorentz factor $\Gamma\approx a\gg 1$.
@@
The problem qualitatively changes when the background plasma is magnetized (B22). Now, in addition to the oscillating electromagnetic field of the wave, there is a background magnetic field $\boldsymbol{B}_{\rm bg}$, which arrests the particle motion across $\boldsymbol{B}_{\rm bg}$ and changes the frame where the particles have zero average momentum (the “fluid frame”). In the media surrounding FRB sources, $B_{\rm bg}$ is well below the radio wave amplitude $E_{0}$ at all radii $r>R_{\rm LC}$ considered in this paper. At the same time, $B_{\rm bg}$ far exceeds a minimum value sufficient to “magnetize” the plasma: the particle Larmor period is much smaller than the duration of the wave packet $T\sim 1$ ms. Besides the wave strength parameter $a$, the problem of wave-plasma interaction has two more dimensionless parameters: the plasma magnetization $\sigma$ and the ratio of gyrofrequency $\omega_{B}=eB_{\rm bg}/mc$ to the wave frequency $\omega$.
The ratio $\omega_{B}/\omega$ plays a particularly important role. B22 showed that waves with $\omega>\omega_{B}$ drive a regular particle motion in the fluid frame: fast oscillations with wave frequency $\omega$ and Lorentz factor $\gamma\approx a$, superposed with (subdominant) slow gyration in the background magnetic field. If $1<\omega_{B}/\omega<a$, gyration is accompanied by a resonant wave-particle interaction each Larmor period, leading to quick stochastic particle acceleration to $\gamma\gg a$. Oscillations in the wave then become energetically subdominant compared to the Larmor rotation with the stochastically pumped $\gamma$. The stochastic heating gives a well-defined mean expectation for the particle energy, which has been found analytically and confirmed numerically.
In addition to heating the plasma, the radio wave also deposits momentum, pushing the plasma along the wave direction. The effective inertia of a plasma with $\sigma\gg 1$ is increased by its magnetic field, which reduces its bulk acceleration. The bulk acceleration leads to compression of the plasma and its magnetic field (Beloborodov, 2021).
@@
Dynamics of a plasma exposed to the radio wave is easiest to investigate in the plane geometry, assuming a uniform magnetized background ahead of the wave. Details of the radio wave are not important, however one needs to choose a concrete shape of the wave packet when simulating the plasma dynamics numerically.
2.2 Wave packet
As an example, consider a nearly harmonic plane electromagnetic wave with amplitude $E_{0}$, frequency $\omega$, and duration $T\gg 2\pi/\omega$. Its strength parameter (defined in Equation (1)) is
$a=\frac{a_{0}}{\sqrt{2}},\qquad a_{0}=\frac{eE_{0}}{mc\omega}.$ (3)
@@ 2.2 Wave packet
The wave speed is weakly affected by the presence of a tenuous plasma, so it propagates with speed of light $c$. It will be convenient to use the coordinate
@@
where the $z$-axis is chosen along the wave propagation direction. The radio wave packet is static in the $\xi$-coordinate and occupies the region $0<\xi<T$. The wave amplitude is zero at the leading edge of the packet $\xi=0$, rises on a timescale $\xi_{\rm rise}<T$, and then drops back to zero at $\xi=T$. In numerical examples we will use packets with the envelope
@@
The cold unperturbed medium ahead of the packet (at $\xi<0$) will be called “upstream.” The upstream medium is described by a uniform mass density $\rho_{\rm u}$ and a uniform magnetic field $\boldsymbol{B}_{\rm u}$. We will focus on the problem with $\boldsymbol{B}_{\rm u}$ oriented perpendicular to the wave propagation direction (as expected in winds, which carry magnetic fields transverse to the radial direction).
@@
2.3 Plasma drift through the wave packet
@@
Here $\overline{\boldsymbol{E}}$ and $\overline{\boldsymbol{B}}$ are the MHD fields; the bar signifies averaging over a timescale longer than the Larmor time (and longer than the wave oscillation period). The average fields can also be denoted as $\boldsymbol{E}_{\rm bg}$ and $\boldsymbol{B}_{\rm bg}$ because they serve as the effective background fields in the problem of particle motion in the wave, as described below.
A natural choice for a reference frame in the wave-plasma interaction problem is such that the upstream plasma is at rest. When discussing wave propagation in a wind, we will also view the problem in the lab frame where the upstream moves with a large speed $\beta_{\rm u}>0$ along the wave propagation direction. In addition, frames with opposite motion, where the upstream plasma has $\beta_{\rm u}<0$, may be useful in numerical simulations. In all these frames, $\boldsymbol{v}_{\rm D}$ is parallel to the $z$-axis, both outside and inside the wave packet. The fluid four-velocity $u_{\rm D}^{\mu}$ then has two non-zero components:
@@
Plasma with any speed $v_{\rm D}=dz/dt$ changes its $\xi$-coordinate with rate
@@
Thus, the MHD fluid always moves toward increasing $\xi$ through the wave packet. When viewed in the rest frame of the upstream medium, the initially static fluid enters the wave packet at $\xi=0$ and develops speed $\beta_{\rm D}\neq 0$ at $\xi>0$. We wish to find the fluid motion in the packet, $\beta_{\rm D}(\xi)$, and the fluid compression factor $C(\xi)=\rho(\xi)/\rho_{\rm u}$.
3 Plasma heating by the wave

### v2

@@ Abstract
When a fast radio burst (FRB) expands from its source through a surrounding tenuous plasma, it strongly heats and compresses the plasma at radii up to $\sim 10^{14}$ cm. The likely central engines of FRBs are magnetars, and their ambient plasma at radii $r\gg 10^{10}$ cm is a magnetized $e^{\pm}$ wind. We formulate basic equations of the FRB-plasma interaction, solve them numerically, and describe the physical picture of the interaction. We find the following: (1) FRBs emitted at $r<r_{\rm stoch}\sim 10^{12}$ cm induce fast stochastic heating and strong compression of the wind, sweeping it like a broom. The outcome of this interaction is determined by the energy losses of the radio wave; we evaluate the parameter space where FRBs survive and escape. (2) At radii $r>r_{\rm stoch}$, FRB induces regular particle oscillations in the radio wave with the standard strength parameter $a$, and drives a compression wave in the wind. At $r>r_{\star}\sim 10^{13}\,$cm, the compression wave becomes locally quasisteady, with compression factor $1+a^{2}$. FRBs avoid damping if they are released into the wind medium outside $r_{\rm damp}\sim 10^{11}$ cm.
@@
I Introduction
Observations of fast radio bursts (FRBs) demonstrate the existence of ultrastrong electromagnetic waves in the Universe. They have characteristic frequencies $\nu=\omega/2\pi\sim 1$ GHz and durations $T\sim 1$ ms (Petroff et al., 2019), and their dimensionless strength parameter is [^1] We use the standard notation $Q_{n}$ for a quantity $Q$ normalized to $10^{n}$ in CGS units, e.g. $\nu_{9}=\nu/10^{9}{\rm Hz}$ and $r_{12}=r/10^{12}{\rm cm}$.
@@ I Introduction
where $m$ is the electron mass, $c$ is the speed of light, $L=cr^{2}\langle E_{\rm wave}^{2}\rangle$ is the isotropic equivalent of wave luminosity; $\langle E_{\rm wave}^{2}\rangle^{1/2}$ is the root-mean-square of the wave electric field when the wave expands to a radius $r$.
FRBs have to propagate through the ambient plasma surrounding the source. The likely central engines of FRBs are neutron stars, which are surrounded by electron-positron plasma. The plasma is strongly magnetized: its magnetic stress $B^{2}/4\pi$ far exceeds the rest-mass energy density $\rho c^{2}$,
@@
where $B$ and $\rho$ are measured in the plasma rest frame. Near the neutron star, the plasma is locked in the closed magnetosphere, rotating with a period $P\sim 1-10$ s. Outside the light cylinder $R_{\rm LC}=cP/2\pi\approx 5\times 10^{9}(P/1{\rm\,s})$ cm, a relativistic wind flows from the magnetosphere, powered by the star’s rotation. Both the magnetosphere and the wind have $\sigma\gg 1$ (see Cerutti & Beloborodov (2017) for a review).
The plasma around neutron stars is tenuous, however it can have a strong effect on FRB propagation. The radio wave with $a\gg 1$ accelerates plasma particles to high energies, and this process can damp the wave, in particular if the energized plasma experiences radiative losses. Strong radiative damping has been demonstrated for FRBs emitted near a magnetar and propagating through its equilibrium outer magnetosphere at radii $r\sim 10^{8}-10^{9}$ cm (Beloborodov, 2021, 2024). In addition, strong radio waves are damped through excitation of small-scale Alfvén waves in the magnetosphere (Golbraikh & Lyubarsky, 2023).
@@
The present paper investigates the FRB-wind interaction using a new method. The problem is stated in section II and then solved in two steps:
First, we describe plasma response to an FRB on a “micro” level by tracking the motion of individual particles (section III). The response can be regular oscillation or stochastic heating (Beloborodov 2022; hereafter B22). It is best viewed in the local center-of-momentum frame of the plasma particles (“fluid frame”). Remarkably, the microscopic response can be described independently of the fluid motion. This fact is illustrated with simulations of particles driven by FRBs in static and accelerating magnetized fluids. The simulations demonstrate the regular and stochastic response, and the transition between the two regimes.
@@
Then, we examine the macroscopic fluid dynamics of plasma exposed to FRBs. The magnetized fluid is pushed and compressed by the FRB on timescales much longer than the particle Larmor period and the wave oscillation period, [^2] The acceleration effect may be thought of as a ponderomotive force exerted by the radio wave. In the presence of strong radiative losses it is more suitable to view it as a scattering process. In any case, it may be described from first principles by solving for particle motion in the wave and then examining the fluid dynamics on macro scales, as done in this paper. so it obeys magnetohydrodynamic (MHD) equations. The effective enthalpy of the MHD fluid is determined by particle motions in the fluid frame (i.e. by the microscopic response to the FRB) and controlled by the Lorentz-invariant parameter $a$. One can think of the micro-level response as a local heating process and the macro-level response as bulk acceleration of the fluid, leading to its compression. Equations for fluid dynamics inside a propagating radio wave packet are formulated in section IV.
It is easy to see why plasma heating by FRBs is inevitably accompanied by its bulk acceleration: any energy $\Delta{\cal E}$ received from the radio wave comes with momentum $\Delta{\cal E}/c$ directed along the wave propagation direction (the wave loses momentum $\Delta{\cal E}/c$). Beloborodov (2021) estimated this effect for FRBs propagating through the dipole magnetosphere of a neutron star, where the FRB-plasma interaction is highly radiative and can be described as a scattering process. A more general and rigorous calculation should be formulated as a time-dependent fluid dynamics problem, using MHD equations and allowing weak or strong radiative losses of the heated plasma. This is attempted in the present paper. The weakly radiative regime becomes particularly relevant when the FRB propagates through the wind far outside the closed magnetosphere.
While heating by FRBs implies push and compression, the compression in turn can change the heating process. Note that the MHD fluid is compressed together with its magnetic field. Once the magnetic field exceeds a certain value, regular particle oscillations in the radio wave change into stochastic heating, as shown below. This change boosts the heating rate and leads to even stronger compression, with enhanced radiative losses of the heated particles. The compression fronts driven by FRBs in magnetar winds involve quite dramatic effects, capable of damping the FRB itself as discussed in later sections of this paper.
@@
Before considering FRBs expanding through a radial wind, we examine the compression front in a plane geometry with a uniform background. First, we show that the problem has a simple steady-state solution (section V), which agrees with the results of Sobacchi et al. (2024). Its compression factor $C_{\star}$ is directly related to the particle energy gained from the radio wave and becomes huge for waves with strength parameter $a\gg 1$. Remarkably, $C_{\star}$ is independent of the plasma magnetization $\sigma$. This fact casts doubt on the relevance of the steady-state solution; in particular, it clearly becomes irrelevant in the vacuum limit ($\sigma\rightarrow\infty$), since the radio wave in vacuum propagates through the background magnetic field with no interaction and no compression effect. A complete picture is revealed by the full time-dependent solution (section VI). It shows that the MHD flow pattern in the propagating wave packet is gradually attracted to the steady state, and this relaxation takes a significant time $t_{\star}$ proportional to $\sigma$. For propagation times $t<t_{\star}$, the steady state with large $C_{\star}$ is not reached; we evaluate the compression factor $C\ll C_{\star}$ in this situation.
@@
These results allow one to investigate the FRB interaction with a magnetar wind in a broad range of radii. We describe the wind in section VII, and then investigate the compression front induced by the FRB in the wind. We identify two zones, with regular and stochastic heating, and find the transition radius $r_{\rm stoch}$ between them. Compression fronts from FRBs emitted in each zone are described in sections VIII and IX. FRB damping by its interaction with the wind is investigated in section X. Our conclusions are summarized in section XI.
II Problem formulation
II.1 The role of magnetic field
@@
For an unmagnetized plasma, particle acceleration in a strong electromagnetic wave is a textbook problem (Landau & Lifshitz, 1975) well known in the context of laser experiments (Bulanov et al., 2015). It was also discussed in the context of early pulsar models (Gunn & Ostriker, 1971). The wave accelerates initially static particles to the Lorentz factor $\gamma=1+a^{2}/2$. The particle oscillating in the wave has a large mean velocity aligned with the wave propagation direction. This mean velocity gives the plasma bulk Lorentz factor $\Gamma\approx a\gg 1$.
@@
The problem qualitatively changes when the background plasma is magnetized (B22). Now, in addition to the oscillating electromagnetic field of the wave, there is a background magnetic field $\boldsymbol{B}_{\rm bg}$, which arrests the particle motion across $\boldsymbol{B}_{\rm bg}$ and changes the frame where the particles have zero average momentum (the fluid frame). In the media surrounding FRB sources, $B_{\rm bg}$ is well below the radio wave amplitude $E_{0}$ at radii $r>R_{\rm LC}$ considered in this paper. At the same time, $B_{\rm bg}$ is sufficiently strong to “magnetize” the plasma: the particle Larmor period is far smaller than the duration of the radio wave packet $T\sim 1$ ms. Larmor rotation couples the plasma to $B_{\rm bg}$, and together they form an MHD fluid, which becomes heated and pushed by the radio wave.
Besides the wave strength parameter $a$, the problem of FRB-plasma interaction has two more dimensionless parameters: the background plasma magnetization $\sigma$ and the ratio of gyrofrequency $\omega_{B}=eB_{\rm bg}/mc$ to the wave frequency $\omega$. The parameters $a$ and $\omega_{B}/\omega$ determine whether the plasma particles execute regular oscillations or experience stochastic heating in the radio wave. In addition to heating, the radio wave also deposits momentum, pushing the plasma along the wave direction. This effect depends on $\sigma$: $\sigma\gg 1$ implies that the effective bulk inertia of MHD fluid is significantly increased by its magnetic field, which can slow down its bulk acceleration. However, the acceleration effect may still be strong in waves with $a\gg 1$.
@@
Dynamics of a magnetized plasma exposed to the radio wave is easiest to calculate in the plane geometry, assuming a uniform magnetized background ahead of the wave packet. Details of the radio wave are not important; however, when simulating the plasma dynamics numerically, one needs to choose a concrete shape of the wave packet.
II.2 Radio wave packet
@@
The wave speed is weakly affected by the presence of a tenuous plasma, and it will be taken equal to $c$ (this approximation is discussed in Appendix A). It is convenient to use the coordinate
@@
where the $z$-axis is along the wave propagation direction. The radio wave packet is static in the $\xi$-coordinate and occupies a finite region $0<\xi<T$. As an example, consider a nearly harmonic plane electromagnetic wave with amplitude $E_{0}$, frequency $\omega$, and duration $T\gg 2\pi/\omega$. The wave electric field,
$E_{\rm wave}=E_{0}(\xi)\sin(\omega\xi),$ (4)
oscillates with a high $\omega$ and a gradually varying amplitude $E_{0}(\xi)$; $E_{0}=0$ at the leading edge of the packet $\xi=0$, rises on a timescale $\xi_{\rm rise}\gg\omega^{-1}$, and then drops back to zero at $\xi=T$.
Using $cE_{\rm wave}=-dA_{\rm wave}/d\xi$ one finds the wave potential at $\xi\omega\gg 1$: $A_{\rm wave}\approx A_{0}\cos(\omega\xi)$ where $A_{0}=cE_{0}/\omega$. So,
$a_{\rm wave}=a_{0}(\xi)\cos(\omega\xi),\qquad a_{0}=\frac{eE_{0}}{mc\omega},$ (5)
where $a_{\rm wave}\equiv eA_{\rm wave}/mc^{2}$. The wave strength parameter is
$a\equiv\langle a_{\rm wave}^{2}\rangle^{1/2}=\frac{a_{0}}{\sqrt{2}}.$ (6)
In numerical examples we will use packets with the envelope
@@
The cold unperturbed medium ahead of the packet (at $\xi<0$) will be called “upstream.” The upstream medium is described by a uniform mass density $\rho_{\rm u}$ and a uniform magnetic field $\boldsymbol{B}_{\rm u}$. We will focus on the case with $\boldsymbol{B}_{\rm u}$ oriented perpendicular to the wave propagation direction (as expected in winds, which carry magnetic fields transverse to the radial direction). [^3] The problem formulation with transverse $\boldsymbol{B}_{\rm bg}$ is not special or degenerate in any way. One could also consider $\boldsymbol{B}_{\rm bg}$ with arbitrary angle $\alpha$ with respect to the wave propagation direction. The general case with any $\alpha\neq 0$ is reduced to the propagation problem with $\alpha=\pi/2$ by using a new reference frame that moves along $\boldsymbol{B}_{\rm bg}$ with Lorentz factor $\gamma=(\sin\alpha)^{-1}$ (Beloborodov, 2024).
@@
II.3 Plasma drift through the wave packet
@@
Here $\overline{\boldsymbol{E}}$ and $\overline{\boldsymbol{B}}$ are the MHD fields; the bar signifies averaging over a timescale longer than the Larmor time (and longer than the wave oscillation period). The average fields can also be denoted as $\boldsymbol{E}_{\rm bg}$ and $\boldsymbol{B}_{\rm bg}$ because they serve as the effective background fields in the problem of test particle motion in the wave, as described below.
A natural choice for a reference frame in the wave-plasma interaction problem is such that the upstream plasma is at rest. When discussing wave propagation in a wind, we will also view the problem in the lab frame where the upstream plasma moves with a large speed $\beta_{\rm u}>0$ along the wave propagation direction. In addition, frames with opposite motion, where the upstream plasma has $\beta_{\rm u}<0$, may be useful in numerical simulations. In all these frames, $\boldsymbol{v}_{\rm D}$ is parallel to the $z$-axis, both outside and inside the wave packet. The fluid four-velocity $u_{\rm D}^{\mu}$ then has two non-zero components:
@@
One can also describe fluid motion using the quantity
This simplifies transformations between different frames moving along $z$. Relativistic addition of velocities corresponds to multiplication of $\kappa$, and changing $\beta_{\rm D}\rightarrow-\beta_{\rm D}$ corresponds to $\kappa\rightarrow\kappa^{-1}$. Doppler transformation of the wave frequency satisfies $\omega/\kappa={\rm inv}$. Transformation of other quantities follow from $a={\rm inv}$ (strength parameter of the wave) and $\omega\xi={\rm inv}$ (phase of the wave). In particular, transformations of the wave electric field $E_{\rm wave}$, wave energy density $U$, and coordinate $\xi$ measured in two reference frames $\cal K$} and $\tilde{\cal K}$ are
$\frac{\tilde{E}_{\rm wave}}{E_{\rm wave}}=\frac{\xi}{\tilde{\xi}}=\frac{\tilde{\omega}}{\omega}=\frac{\tilde{\kappa}}{\kappa},\qquad\frac{\tilde{U}}{U}=\frac{\tilde{\kappa}^{2}}{\kappa^{2}}.$ (11)
Hereafter, we choose $\tilde{\cal K}$ as the local rest frame of the fluid, [^4] Throughout this paper, we use tilde for quantities measured in the fluid frame $\tilde{\cal K}$. Tilde will be omitted if this does not lead to confusion: we omit tilde for thermodynamic quantities (e.g. pressure $P$ and enthalpy $w$, which are defined only in frame $\tilde{\cal K}$). which corresponds to
$\tilde{\beta}_{\rm D}=0,\qquad\tilde{\kappa}=1.$ (12)
Plasma with any speed $v_{\rm D}=dz/dt$ (in any frame) changes its $\xi$-coordinate with rate
@@
Thus, MHD fluid always moves toward increasing $\xi$ through the wave packet. When viewed in the rest frame of the upstream medium, the initially static fluid enters the wave packet at $\xi=0$ and develops speed $\beta_{\rm D}\neq 0$ at $\xi>0$. We wish to find the fluid motion in the packet, $\beta_{\rm D}(\xi)$, and the fluid compression factor $C(\xi)=\rho(\xi)/\rho_{\rm u}$.
III Plasma heating by the radio wave
The response of a magnetized plasma to a propagating radio wave packet can be disentangled if we first focus on the “micro level” and consider the motion of individual particles in the local fluid frame $\tilde{\cal K}$. In this frame, the fluid is at rest and the wave-plasma interaction may be viewed as a local heating effect: the strong high-frequency wave energizes particles to Lorentz factors $\tilde{\gamma}\gg 1$. This effect was calculated in B22 by tracking the motion of test particles. The heating depends on the wave strength parameter $a$. For smooth packets, such that $a$ varies on a scale $\Delta\tilde{\xi}\gg a\tilde{\omega}_{B}$ (typically satisfied by FRBs), the results for $\tilde{\gamma}$ may be summarized as follows (see also Appendix B).
(1) Waves with frequency $\tilde{\omega}$ exceeding the gyrofrequency $\tilde{\omega}_{B}=e\tilde{B}_{\rm bg}/mc$ drive regular particle oscillations with the average Lorentz factor
$\langle\tilde{\gamma}\rangle=\sqrt{1+a^{2}}.$ (14)
The local $\langle\tilde{\gamma}\rangle$ in the packet simply tracks the local strength parameter of the wave $a$.